OpenStudy (anonymous):

Slope-intercept help! Medal/Follow If possible write the following equation in slope-intercept form. y=−4 a) What is the value of m when written in slope-intercept form? If the equation cannot be written in slope-intercept form, enter none in the box. m= b) What is the value of b when written in slope-intercept form? If the equation cannot be written in slope-intercept form, enter none in the box. b=

OpenStudy (anonymous):

y = -4

OpenStudy (anonymous):

Nevermind, I figured it out because I am a genius

OpenStudy (anonymous):

so proud

OpenStudy (anonymous):

@juliannescriv lol, can you help me with another one please?

OpenStudy (anonymous):

yepyep

OpenStudy (anonymous):

If possible write the following equation in slope-intercept form. 3x+4y+26=x−y a) What is the value of m when written in slope-intercept form? If the equation cannot be written in slope-intercept form, enter none in the box. m= b) What is the value of b when written in slope-intercept form? If the equation cannot be written in slope-intercept form, enter none in the box. b=

OpenStudy (anonymous):

those question marks are x-y

OpenStudy (anonymous):

-.4x and 5.2 are what i got but i dont think they are right

OpenStudy (anonymous):

2x+4y+26=-y 2x+26=-5y -2/5x-26/5=y

OpenStudy (anonymous):

the -.4x is right, but the 5.2 needs to be negative too

OpenStudy (anonymous):

5.2 is right, first one is wrong somehow :(

OpenStudy (anonymous):

this is the message i got: Your answer isn't a number (it looks like a formula that returns a number)

OpenStudy (anonymous):

because it wouldnt be x

OpenStudy (anonymous):

did you put -.4 or -.4x

OpenStudy (anonymous):

ohhhhh hahaha thank you so much! :)

OpenStudy (anonymous):

mhm :)